OpenStudy (muddahamburglar):

(where does the 1/2 come from in this) Simplify cos(a+pi/3) cos a cos π/3 - sin a sin π/3 cos a [ 1/2 ] - (√3/2) sin a (1/2) [ cos a - √3 sin a ]

OpenStudy (john_es):

The 1/2 come from the fact that: \[\cos(\pi/3)=1/2\]

OpenStudy (john_es):

And\[\sin(\pi/3)=\sqrt{3}/2\]

OpenStudy (muddahamburglar):

@John_ES in the last step where does the two go from sqtrt3/2?

OpenStudy (muddahamburglar):

oh wait nvm i typed it out wrong that's why it was gone lol

OpenStudy (freckles):

@muddahamburglar are you still having issues with this question?

OpenStudy (muddahamburglar):

yes i am @freckles

OpenStudy (freckles):

can you tell me which part?

OpenStudy (phi):

notice in (1/2) [ cos a - √3 sin a ] the ½ is "outside" the brackets which act as parens. if you distribute the ½ you will get back ½ cos a - sqr(3)/2 sin a you can "factor out" the ½ and go back to (1/2) [ cos a - √3 sin a ]

OpenStudy (muddahamburglar):

@freckles actually i think i get now thank you so much though!

OpenStudy (freckles):

k